# Strong Converse Exponent of Quantum Dichotomies

### v1

@@ VI Discussion
In this work, we determine the exact strong converse exponent of quantum dichotomies based on the purified distance. It is characterized by the sandwiched Rényi divergence $D_{\beta}(\rho_{1}\|\sigma_{1})$ of the initial quantum dichotomy with $\beta\in(1,\infty)$ and the sandwiched Rényi divergence $D_{\alpha}(\rho_{2}\|\sigma_{2})$ of the target quantum dichotomy $(\rho_{2},\sigma_{2})$ with $\alpha\in(\frac{1}{2},1)$. Hence, we find a new type of multi-variate quantum Rényi information measure involving four quantum states — with an exact operational meaning. The combination of this result and an improved Fuchs–van de Graaf inequality also directly leads to the strong converse exponent based on the trace distance when $\rho_{2}$ is a pure state. We further believe that this formula still holds when $\rho_{2}$ is a general quantum state, but leave it as an open question. While recent results have established lower bounds for the strong converse exponents under the trace distance for some tasks Wilde2022distinguishability ; SalzmannDatta2022total ; SGC2022optimal , deriving a matching upper bound seems a difficult problem. The improved Fuchs–van de Graaf inequality provides a new idea to solve this kind of problem when dealing with pure states. In this case, the strong converse exponent under the trace distance can be transformed into that under the purified distance. For determining the strong converse exponent under the purified distance, there are then already a few techniques developed MosonyiOgawa2017strong ; LiYao2024operational .
@@ Acknowledgments
MB and YY acknowledge funding by the European Research Council (ERC Grant Agreement No. 948139) and MB acknowledges support from the Excellence Cluster - Matter and Light for Quantum Computing (ML4Q).
VII Appendix
@@
Lemma 12
For any pure state $\varphi$ and subnormalized state $\rho$, we have
$P(\rho,\varphi)\leq\sqrt{d(\varphi,\rho)}.$ (48)
Proof. By direct calculation, we have
$d(\varphi,\rho)=\operatorname{Tr}(\varphi-\rho)_{+}\geq\operatorname{Tr}(\varphi-\rho)\varphi=1-\operatorname{Tr}\rho\varphi=1-F^{2}(\rho,\varphi)=P^{2}(\rho,\varphi).$ (49)
[…]
Lemma 13
@@
Proof. By Ulhmann’s theorem Uhlmann1976transition , there exists purified states $\phi_{\rho}$, $\phi_{\sigma}$ and $\phi_{\tau}$ of $\rho$, $\sigma$ and $\tau$ such that
@@
By direct calculation, we have
$\begin{split}F^{2}(\tau,\sigma)&\geq F^{2}(\phi_{\tau},\phi_{\sigma})\\ &=\operatorname{Tr}\phi_{\tau}\phi_{\sigma}\\ &=\operatorname{Tr}\phi_{\sigma}\phi_{\rho}+\operatorname{Tr}\phi_{\sigma}(\phi_{\tau}-\phi_{\rho})\\ &\geq\operatorname{Tr}\phi_{\sigma}\phi_{\rho}-\operatorname{Tr}(\phi_{\rho}-\phi_{\tau})_{+}\\ &=F^{2}(\rho,\sigma)-P(\phi_{\rho},\phi_{\tau})\\ &=F^{2}(\rho,\sigma)-P(\rho,\tau)\\ &\geq F^{2}(\rho,\sigma)-\sqrt{\|\rho-\tau\|_{1}}.\end{split}$ (51)
[…]
The following three lemmas are used in our proofs and they have been proved in LiYao2024operational .
Lemma 14
Let $\rho,\sigma\in\mathcal{S}(\mathcal{H})$, and let $\mathcal{H}=\bigoplus_{i\in\mathcal{I}}\mathcal{H}_{i}$ decompose into a set of mutually orthogonal subspaces $\{\mathcal{H}_{i}\}_{i\in\mathcal{I}}$. Suppose that $\sigma=\sum\limits_{i\in\mathcal{I}}\sigma_{i}$ with ${\operatorname{supp}}(\sigma_{i})\subseteq\mathcal{H}_{i}$. Then
@@
Lemma 16
Let $\rho,\sigma\in\mathcal{S}(\mathcal{H})$ and $\tau\in\mathcal{P}(\mathcal{H})$, and suppose ${\operatorname{supp}}(\sigma)\not\perp{\operatorname{supp}}(\tau)$. Fix $\alpha\in(\frac{1}{2},1)$ and $\beta\in(1,+\infty)$ such that $\frac{1}{\alpha}+\frac{1}{\beta}=2$. Then
$\frac{2\alpha}{1-\alpha}\log F(\rho,\sigma)\leq D^{*}_{\beta}(\rho\|\tau)-D^{*}_{\alpha}(\sigma\|\tau).$ (54)

### v2

@@ V Discussion
In this work, we determine the exact strong converse exponent of quantum dichotomies based on the purified distance. The combination of this result and an improved Fuchs–van de Graaf inequality also directly leads to the strong converse exponent based on the trace distance when $\rho_{2}$ is a pure state. We further believe that this formula still holds when $\rho_{2}$ is a general quantum state, but leave it as an open question. While recent results have established lower bounds for the strong converse exponents under the trace distance for some tasks [8, 37, 38], deriving a matching upper bound seems a difficult problem. The improved Fuchs–van de Graaf inequality provides a new idea to solve this kind of problem when dealing with pure states. In this case, the strong converse exponent under the trace distance can be transformed into that under the purified distance. For determining the strong converse exponent under the purified distance, there are then already a few techniques developed [18, 19].
@@ Acknowledgments
The authors would like to thank Mark Wilde for comments. MB and YY acknowledge funding by the European Research Council (ERC Grant Agreement No. 948139) and MB acknowledges support from the Excellence Cluster - Matter and Light for Quantum Computing (ML4Q).
VI Appendix
@@
Lemma 10.
For any pure state $\varphi$ and quantum state $\rho$, we have
$P(\rho,\varphi)\leq\sqrt{d(\rho,\varphi)}.$ (47)
Proof: By direct calculation, we have
$d(\rho,\varphi)=\operatorname{Tr}(\varphi-\rho)_{+}\geq\operatorname{Tr}(\varphi-\rho)\varphi=1-\operatorname{Tr}\rho\varphi=1-F^{2}(\rho,\varphi)=P^{2}(\rho,\varphi).$ (48)
We are done.
Lemma 11.
@@
Proof: By Ulhmann’s theorem [39], there exists purified states $\phi_{\rho}$, $\phi_{\sigma}$ and $\phi_{\tau}$ of $\rho$, $\sigma$ and $\tau$ such that
@@
It is easy to see that
$\begin{split}F^{2}(\tau,\sigma)&\geq F^{2}(\phi_{\tau},\phi_{\sigma})\\ &=\operatorname{Tr}\phi_{\tau}\phi_{\sigma}\\ &\geq\operatorname{Tr}\phi_{\rho}\phi_{\sigma}-\operatorname{Tr}(\phi_{\rho}-\phi_{\tau})_{+}\\ &=F^{2}(\rho,\sigma)-P(\phi_{\rho},\phi_{\tau})\\ &=F^{2}(\rho,\sigma)-P(\rho,\tau)\\ &\geq F^{2}(\rho,\sigma)-\sqrt{\|\rho-\tau\|_{1}}.\end{split}$ (50)
The following three lemmas are used in our proofs and they have been proved in [19, Lemma 30] and [19, Lemma 31], respectively.
Lemma 12.
Let $\rho,\sigma\in\mathcal{S}(\mathcal{H})$, and let $\mathcal{H}=\bigoplus_{i\in\mathcal{I}}\mathcal{H}_{i}$ decompose into a set of mutually orthogonal subspaces $\{\mathcal{H}_{i}\}_{i\in\mathcal{I}}$. Suppose that $\sigma=\sum\limits_{i\in\mathcal{I}}\sigma_{i}$ with $\mathop{\rm supp}(\sigma_{i})\subseteq\mathcal{H}_{i}$. Then